1 - The return on shares of Valley Transporter is predicted under the following various economic conditions:
Recession -0.12
Normal +0.06
Boom +0.24
If each economy state has the same probability of occurring, what is the variance of the stock?

Place your answer in decimal form using four decimal places.

2. The return on shares of the Orange Company are predicted under the following states of nature. The states of nature are all equally likely, and because there are a total of three states, each state has a 33.333% chance of occurring.

Recession -0.14
Normal +0.04
Boom +0.23

What is the standard deviation of Orange?

* Place your answer in decimal form, for example as say .0675 and not 6.75.

3. Toyota Corp.'s stock is $30 per share. Its expected return is 24% and variance is 13%. Honda Corp.'s stock is $22 per share. Its expected return is 15% and variance is 8%. Benz Corp.'s stock is $45 per share. Its expected return is 14% and variance 4%. What would be the expected return of a portfolio consisting of 50% Toyota and 50% Honda?
%

* Place your answer in percentage form, say 5.99% and not .0599

4.Toyota has an expected return of 21%, and a variance of 0.010. Honda has an expected return of 18%, and a variance of 0.005. The covariance between Toyota and Honda is 0.08. Using these data, calculate the variance of a portfolio consisting of 50% Toyota and 50% Honda.

please help me
Answer #1

Answer to Question 1:

Probability of each economy = 1/3

Expected return = (1/3) * (-0.12) + (1/3) * 0.06 + (1/3) * 0.24
Expected return = 0.06

Variance = (1/3) * (-0.12 - 0.06)^2 + (1/3) * (0.06 - 0.06)^2 + (1/3) * (0.24 - 0.06)^2
Variance = 0.0216

Answer to Question 2:

Expected return = (1/3) * (-0.14) + (1/3) * 0.04 + (1/3) * 0.23
Expected return = 0.0433

Variance = (1/3) * (-0.14 - 0.0433)^2 + (1/3) * (0.04 - 0.0433)^2 + (1/3) * (0.23 - 0.0433)^2
Variance = 0.022822

Standard Deviation = (0.022822)^(1/2)
Standard Deviation = 0.1511